OpenStudy (anonymous):

log4 x + log4(x + 4) = 0 find real number solutions

OpenStudy (anonymous):

Do you know you log rules?

OpenStudy (anonymous):

i know some but for my hw we dont have explanations for ones like htis

OpenStudy (lgbasallote):

use product law of logs..change to exponential..solve

OpenStudy (anonymous):

i dony know how to do this problem

OpenStudy (anonymous):

Things will work out for the best.

OpenStudy (anonymous):

log4 x + log4(x + 4) = 0 log4 [x ( x +4) ] = 0 [x ( x +4) ] = 4^0 = 1 x^2 + 4x - 1 = 0 x = -2 +/- sqrt (5)

OpenStudy (anonymous):

Just plug into your calculator to get result!